X-rays of wavelengths 1×10-10 are scattered from a carbon block to undergo a 90° Campton scattering.

a) calculate the Campton shift

b) what is the kinetic energy of the recoiling electron


1
Expert's answer
2021-12-20T10:29:56-0500

The shift is


"\lambda-\lambda'=\frac h{m_ec}(1-\cos\theta)=2.45·10^{-12}\text{ m}."

The kinetic energy of the recoiling electron:


"E_r=\frac{hc/\lambda}{1+h/(m_e\lambda c)(1-\cos\theta)}=12.1\text{ keV}."